OpenStudy (ammarah):

For the sequence, describe the pattern, write the next term, and write a rule for the nth term 2,4,8,16

hartnn (hartnn):

did you notice any pattern ?

OpenStudy (ammarah):

x2

hartnn (hartnn):

\(x^2\) are you sure ? or you mean \(2^x\) ?

OpenStudy (ammarah):

i mean it times 2

hartnn (hartnn):

ohh, then you're correct.

hartnn (hartnn):

so what term comes after 16 ?

OpenStudy (ammarah):

32

OpenStudy (anonymous):

help me please hartnn from bigpapa021201

hartnn (hartnn):

yes. so, for n=1, you have 2 = 2^1 for n=2, you have 4 = 2^2 and so on.... can you find the n'th term rule now ?

OpenStudy (ammarah):

wait so do i write the nth term rule for 32?

OpenStudy (anonymous):

help me ammarahplease from bigpapa021201

OpenStudy (anonymous):

so one help me

hartnn (hartnn):

n'th term rule is for entire sequence

OpenStudy (ammarah):

so i would do it for each one?

hartnn (hartnn):

finding n'th term is just once, per sequence, and that n'th term rule applies to all terms,... let me give you example 3,6,9,12.... so, the rule is n times 3, so, f(n) here will be 3n what about 2,4,8,16,32... try ?

OpenStudy (ammarah):

2n

hartnn (hartnn):

not actually, the 2n sequence is 2,4,6,8,10,12....is that your sequence ?

hartnn (hartnn):

note that the terms increase exponentially

OpenStudy (ammarah):

2^n

hartnn (hartnn):

thats absolutely correct! :)

OpenStudy (ammarah):

how about 3,5,7,9 would this be 2+n

hartnn (hartnn):

2nd term , put n= 2 in 2+n so, is the 2nd term = 2+2 = 4 ?? try again ...

hartnn (hartnn):

you're close though.... think of the sequence, 2,4,6,8,10.....and that you add 1 to each term of these

OpenStudy (ammarah):

? idk

hartnn (hartnn):

learnt about arithmetic sequence and common difference yet ?

OpenStudy (ammarah):

yes

hartnn (hartnn):

whats the common difference of this sequence ?

hartnn (hartnn):

\(a_n = a+ (n-1)d\) here a= first term = 3 d= ...?

OpenStudy (ammarah):

2

hartnn (hartnn):

yes, so the rule here will be an = 3+ (n-1)*2 simplify this!

OpenStudy (ammarah):

how about this one: 1/1,1/4,1/9,1/16

hartnn (hartnn):

you need n'th term ?? look at denominators... 1,4,9,16... notice any pattern ?

OpenStudy (ammarah):

is the same yes write a rule for the nth term

OpenStudy (ammarah):

i dont see a pattern

hartnn (hartnn):

hint : squares of integers

OpenStudy (ammarah):

2^n

OpenStudy (ammarah):

so how would i write that

hartnn (hartnn):

not actually, the denominator pattern is n^2 n= 1, >>>>1 n=2 >>>>>2^2 = 4 n=3 >>>>>3^2 =9 ... got this ?

OpenStudy (ammarah):

yes

hartnn (hartnn):

and the numerator is always 1 so, your n'th term rule is simply 1/n^2 thats it!

OpenStudy (ammarah):

how would i write the series using summation notation 2/3+4/5+6/6+8/7

hartnn (hartnn):

numerator sequence = 2,4,6,8.... i don't see any pattern in denominator....is the question correct ?

OpenStudy (ammarah):

yeah..

OpenStudy (ammarah):

I have to go to sleep...tell me if u figure out the prob please

hartnn (hartnn):

ok, and good night! sweet dreams :)